If you mix 140mL of a 60% acid solution with 200 mL of a 30% solution yu get a 340 mL of a _____% solution.
Found 2 solutions by checkley75, ptaylor:
Answer by checkley75(3666)

.6*140+.3*200=X*340
84+60=340X
144=340X
X=144/340
X=.4235% SOLUTION

Answer by ptaylor(2198)

Let x=the % acid solution in the final 340 ml mixture.
Now we know that the pure acid in the 140 ml solution (0.60)(140) plus the amount of pure acid in the 200 ml solution (0.30)(200) has to equal the pure acid in the final 340 ml solution (340)(x). So our equation to solve is:
0.60(140)+0.30(200)=340x get rid of parens
84+60=340x
144=340x divide both sides by 340
x=.4235 or 42.35%--------------------------------------------ans

CK
0.60(140)+0.30(200)=340(.4235)
84+60=144
144=144

Hope this helps----ptaylor